# Disjoint supercyclicity in Banach algebras

### v7

@@ 3.3. Applications to generalized weighted shifts on function sequence spaces
(where $\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is the function from Lemma 3.8 with compact ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$). By the proof of Lemma 3.9 (see [15]) one can deduce that the condition ($\mathcal{P}$) is satisfied in this case.
Let now $\alpha=\alpha_{1}=\cdots=\alpha_{N}.$ For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\Phi_{n,\ell}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is not hard to check that the system $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is again a system of isometric isomorphisms also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}.$ Moreover, it is disjoint apreiodic and satisfies the condition (R) with respect to
$\{\tilde{p}_{((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}\}.$
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\mathbb{R}),b_{n,\ell}\in\mathcal{A}_{1}$ and $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be as above in the previous case (when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$).
Now, given $\epsilon_{1}\in(0,1),$ we choose some $\epsilon_{3}\in(\epsilon_{1},1).$ Since ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$ and ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is compact, by Lemma 3.8, we can find some function $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\in\mathcal{C}_{\tau\circ\alpha^{j}}$ such that $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}(t)=1$ for all $t\in supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}},\text{ }supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}$ is compact and $supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\subseteq|\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{3}\right]).$ Therefore, for each $n\in\mathbb{N},\ell\in\{1,\dots,N\},J\in\mathbb{N},\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}$ and any collection of compact subsets
$K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})},$
we let $b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1})_{j+r_{\ell}n}=\frac{v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j+r_{\ell}n})},\epsilon_{1}}\bigr),\epsilon_{3}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ (where we choose $\epsilon_{3}\in(\epsilon_{1},1)$) and $(b_{n,\ell,(\epsilon_{2},\epsilon_{1}J)}^{-1})_{j+r_{\ell}n}=0$ else.
By some calculations, it is not hard to deduce the next corollary as a corollary of Theorem 2.1.
Corollary 3.11.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and finite collection
$\{K_{\epsilon_{J}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{-J+1}}^{(\tau\circ\alpha^{-J+1})},\ldots,K_{\epsilon_{0}}^{\tau},\ldots,K_{\epsilon_{J}}^{(\tau\circ\alpha^{J})}\}$
of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$

### v8

@@ 5. Dynamics on weighted spaces of Radon measures
Now we will prove $(2)\Rightarrow(1)$. Given non-empty open subsets $O,O_{1},\dots,O_{N}$ of $M(\Omega)$, by Remark 5.3 we can find some $\mu\in O,$ $\nu_{\ell}\in O_{\ell}(\ell\in\{1,\dots,N\})$, and a compact $K\subseteq O$ such that
$|\mu|(K^{c})=0\quad\text{and}\quad|\nu_{\ell}|(K^{c})=0\quad\text{for each }\ell\in\{1,\dots,N\}.$
Also, there exists some $\delta>0$ such that
$B(\mu,\delta)\subseteq O\quad\text{and}\quad B(\nu_{\ell},\delta)\subseteq O_{\ell}$
for each $\ell\in\{1,\dots,N\}$. Let
$\varepsilon:=\frac{\delta^{2}}{\left(2(N+1)\left(2+\frac{1}{m_{K}}\right)+\delta\right)^{2}}.$
Choose $F\in\mathcal{F}$ and the families $\{A_{z}\}_{z\in F}$, $\{B_{z,1}\}_{z\in F},\dots,\{B_{z,N}\}_{z\in F}$ satisfying the assumptions of (2) with respect to $K$ and $\varepsilon$.
Let $\widetilde{\mu}_{z},\widetilde{\nu}_{z,1},\dots,\widetilde{\nu}_{z,N}$ be the measures in $M(\Omega)$ given by
$\widetilde{\mu}_{z}(E)=\mu(E\cap A_{z}^{c}\cap K)$
and
$\widetilde{\nu}_{z,\ell}(E)=\nu_{\ell}(E\cap B_{z,\ell}^{c}\cap K)$
for every Borel subset $E$ of $\Omega$, and each $z\in F,\ell\in\{1,\dots,N\}$.
Then, for each distinct $s,\ell\in\{1,\dots,N\}$,
$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\rho\,d\big(\widetilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)$
for all $z\in F$ and every Borel subset $E$ of $\Omega$. Hence,
$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E\cap\alpha_{z,\ell}\big(\alpha_{z,s}^{-1}(B_{z,s}^{c}\cap K)\big)}\rho\frac{(w_{z,\ell}\circ\alpha_{z,\ell}^{-1})}{(w_{z,s}\circ\alpha_{z,\ell}^{-1})}\,d\big(\tilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)=$
$\int_{\alpha_{z,s}\!\big(\alpha_{z,\ell}^{-1}(E)\big)\cap\big(B_{z,s}^{c}\cap K\big)}(\rho\circ\alpha_{z,l}\circ\alpha_{z,s}^{-1})\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})}{(w_{z,s}\circ\alpha_{z,s}^{-1})}\,d\tilde{v}_{z,s}$
for every $z\in F,$ each distinct $s,\ell\in\{1,\dots,N\}$ and every Borel subset $E$ of $\Omega$. By exactly the same arguments as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ we can conclude that
$\Big\|\big(T^{*}_{z,\ell}(S^{*}_{z,s}(\widetilde{\nu}_{z,s}))\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\left((\rho\circ\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(t)\right)\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})(t)}{(w_{z,s}\circ\alpha_{z,s}^{-1})(t)}\|\nu_{s}\|.$
Similarly, as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ and [24, Proposition 4.1] part $ii)\Rightarrow i)$, we can deduce that
$\Big\|\big(T^{*}_{z,\ell}(\widetilde{\mu}_{z})\big)^{(\rho)}\Big\|\leq\sup_{t\in A_{z}^{c}\cap K}\big(((\rho\circ\alpha_{z,\ell})(t))\,(w_{z,\ell}(t))\big)\|\mu\|$
and
$\Big\|\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\big((\rho\circ\alpha_{z,s}^{-1})(t)\big)\,\big((w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\big)\|\nu_{s}\|$
for all $z\in F$ and $s,\ell\in\{1,\dots,N\}$. For each $z\in F$, set
$\psi_{z}=\widetilde{\mu}_{z}+\frac{\sqrt{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}{\sqrt{\displaystyle\sum_{l=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\Big)}}\;\sum_{s=1}^{N}S^{*}_{z,s}(\widetilde{\nu}_{z,s}).$
By the choice of $\varepsilon$, the triangle inequality we can deduce that $\psi_{z}\in O$ and
$\quad\sqrt{\frac{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})(t)\Big)}{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}\;T^{*}_{z,s}(\psi_{z})\in V_{s},$
for all $z\in F$ and $s\in\{1,\dots,N\}$. This proves the implication $(2)\Rightarrow(1)$. ∎
Acknowledgement: I am grateful to Professor S. M. Tabatabaie for suggesting linear dynamics on solid spaces as the topic for my research and for introducing to me the relevant literature.